OpenStudy (anonymous):

You have a 2.00 kg block of lead. Lead melts at 327.5°C. CPb = 130.0 J/kg°C, and (delta)Hf for lead is 2.04 × 10^4 J/kg. Say you start at room temperature (25.0°C). How much heat must you transfer to melt all the lead?

OpenStudy (shamim):

Heat needed to get 327.5 degree celcius Q=m*Cpb*(327.5-25) Q=2*130*(327.5-25=?

OpenStudy (shamim):

Now heat needed to melt the lead Q=m*latent heat of fusion of lead Q=2*2.04*10^4=?

OpenStudy (shamim):

Now add those 2 Q to get total heat needed

OpenStudy (shamim):

Response plz!!!!!

OpenStudy (anonymous):

78650 + 40800 = 119450 ?

OpenStudy (anonymous):

@shamim

OpenStudy (matt101):

@jaysabelle that's right!

OpenStudy (anonymous):

@matt101 Thank you!